OpenStudy (anonymous):

How can I show that x=acost+h, y=bsint+k, 0

OpenStudy (anonymous):

x=acos(t)+h turns into\[\frac{x-h}{a}=\cos t\]do the same with the other eqn. then substitute your values into the unit circle identity\[x ^{2}+y ^{2}=1\]

OpenStudy (anonymous):

oops, I meant \[\cos ^{2}t+\sin ^{2}t=1\]